OpenStudy (anonymous):

log function: 2^x−4+10=22

OpenStudy (anonymous):

I assume it's this\[2^{x-4} + 10 = 22\]and not this \[2^{x}-4+10\] How can you simplify the equation?

OpenStudy (anonymous):

yes it's the first one

OpenStudy (anonymous):

any ideas for first steps?

OpenStudy (anonymous):

I have no idea, that's why I'm asking

OpenStudy (anonymous):

2^x=22-10+4 2^x=16 Now transform this into a log equation.

OpenStudy (anonymous):

The painfully obvious one, using the inverse property of addition and the fact that if b = c, then a+b = a+ c, is\[2^{x-4}+10=22\]\[2^{x-4}+10 + (-10) = 22 + (-10)\]\[2^{x-4}+0 = 12\]\[2^{x-4}=12\]

OpenStudy (anonymous):

If you know how to exponential equations as logarithmic ones, you can do that without much difficulty now.

OpenStudy (anonymous):

Ok, I see. Thank You